What is the kinetic energy. at the maximum highest point if initial kinetic energy be KE ?

Grade:11

1 Answers

Dhruv
6 years ago
If the object is thrown vertically upward, at highest point its velocity becomes zero. Hence kinetic energy is zero.
 
If the object is thrown with initial velocity u at angle theta \Theta, at highest point its velocity will not be zero.
Initial kinetic energy = (1/2) m u2 = E
Final kinetic energy = (1/2) m (u cos \Theta)2 = E (cos2\Theta)